A solid ball has a mass of 200g and a radius of 20cm3 what is the density? 7

Grade:7

1 Answers

Vikas TU
6 years ago
Dear Student,
Volume = 4/3 π x r^3= 4/3x 3.14 x 8000= 33409 cm^3 
Mass is 200 gm. 
Along these lines, thickness is 200/33409=5.9x 10^-3 gm/cc
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)